# Nematic order from phase synchronization of shape oscillations

### v1

@@ Abstract
Unlike solids and simple liquids, orientational order of constituent particles is a defining feature of soft and biological matter [1, 2, 3, 4]. Classic theories predict the emergence of nematic order: Maier-Saupe theory based on anisotropic attractive interactions [5, 6] and Onsager’s theory based on the gain in translational entropy at the expense of lowered orientational entropy of the particles [7]. Athermal, active and external fields [8, 9, 10, 11, 12] can also drive the isotropic - nematic transition. Notably, all these approaches rely on interaction between anisotropic particles. They also ignore a characteristic feature ubiquitous in soft and biological matter, namely the deformability of the constituents e.g., as in polymers, soft colloids, emulsions and cells [13]. In this letter, we show how shape deformability in conjunction with shape anisotropy can induce nematic order of non-interacting particles in a periodically driven system, via a phase synchronization mechanism.
@@
Figure 1: Emergence of nematic order for non-interacting deformable particles in an oscillatory shear flow. (a) Schematic of the system showing deformable particles suspended in a time-dependent shear flow with both a mean (red) and an oscillatory (blue) component. The ially i ropic state of random orientations $\theta_{i}(0)$ and varying shape parameters $r_{i}(0)$ develops into (b) a phase synchronized state at later times. Early and late time kymographs illustrating the temporal evolution of orientation of (c) deformable and (d) rigid particles. (e) The corresponding nematic order parameter $S(t)$ for deformable (black) and rigid (grey) particles. Grey ellipses at top of (c) are a guide for particle orientation while phase synchronized. The plots are for $N=200$, $A=0.5$ and $T=20$.
@@
Figure 3: Transient dynamics in state-space. Snapshots of trajectories in the state-space as points while passing through the $(r,2\theta)$ cross section of the torus at three different times. At $t=0$, random initial conditions (open circles); at $t=T$, trajectories converged to the limit cycle of the steady shear system (yellow filled circles on the red curve); and at $t=14T$, trajectories (black filled circles) converging onto the stable point (green filled circle) that lies on the limit cycle of the oscillatory system (cyan curve). The unstable point (pink cross) corresponds to the unstable limit cycle.
Figure 4: (a) Arnold tongues: regions of phase synchronization induced nematic order (violet) in the $(A,T)$ parameter space. The inset shows the expanded view of the first tongue. (b)-(c) Limit cycles in $(r,2\theta)$ projection of the state-space, corresponding to two points marked in (a): $T=19.5\lesssim T_{n}$ in (b) and $T=20.7\gtrsim T_{n}$ in (c). The color on the limit cycle denotes the state-space speed $\sqrt{(dr/dt)^{2}+(d\theta/dt)^{2}}$. In each figure, the limit cycle for the case of steady shear ($T_{n}\approx 20.02$) is plotted in red. The grey and red ellipses illustrate the aspect ratio of the deformed particle (not to scale) as they pass through the ghost region (near the orientation $\theta=0$) in the oscillatory and in the steady shear flow respectively. Compared to particles in steady shear, in (b) particles reduce their anisotropy as they pass through the ghost region to reduce their period. Conversely, in (c) particles increase their anisotropy to increase their period.
@@
Emergence of nematic order – Simulations show that, over time, all particles tumble in the flow at the same driving frequency $\omega$, thus frequency locking their rotational motion to the imposed oscillatory shear flow. More importantly, we find that the particles also phase lock their rotational motion with the imposed flow (see Sec. IA of [36]). This phase locking of each particle’s orientation with the imposed oscillatory flow implies the synchronization in the motion of all particles with each other as well, and we obtain a state of collective behavior where orientations of all non-interacting deformable particles are the same at all, sufficiently long, times ($t\gg{2\pi}/{\omega}$). Thus, nematic order emerges in the system as indicated schematically in Fig. 1(b) (see Supplemental Video S1 and Sec. IIA of [36]).
Kymographs illustrating the evolution of the orientation of particles $\theta_{i}(t)$ are shown in Fig. 1(c) for early and late times. The isotropy in the initial orientation of particles gradually fades away with time, and at long times the phase-synchronized state emerges with tumbling frequency $\omega$ and no variations in $\theta_{i}$ between different particles. In contrast, the kymograph obtained for rigid, elongated particles, shown in Fig. 1(d) does not exhibit phase locking and the system does not develop nematic order (see Supplemental Video S2 and Sec. IIB of [36]).
@@
This analysis reveals that phase synchronization, and hence the emergence of nematic order, is a direct consequence of the deformability of particles. We now investigate the underlying dynamical mechanism responsible for phase synchronization of the deformable, tumbling particles.
Phase synchronization– Eqs. (1)-(2) constitute a non-autonomous dynamical system due to the time-dependent driving $\dot{\gamma}(t)$, but introducing a new phase variable $\phi_{i}=\phi=\omega t$ or,
@@
results in Eqs. (1)-(3) that form an autonomous $3$D dynamical system for a single particle and allow us to pursue a state-space analysis.
@@
Physical mechanism– Fig. 4(a) shows the range of amplitude, $A$, and oscillatory driving period, $T$, for which phase synchronization occurs. The synchronization regions (violet) occur periodically in bands near multiples of $T_{n}$, the natural time period of tumbling in a steady shear flow ($A=0$). These structures in Fig. 4(a) are the well-known Arnold tongues [14] and they come in and out of existence in parameter space via saddle-node bifurcation of cycles in state space (see Sec. IB of [36]).
@@
The finite width of the tongues can be physically rationalized by analyzing the extent to which a deformable particle can modulate the rate of tumbling and shape changes to match the driving period $T$ of the oscillatory shear. For example, within the primary Arnold tongue in Fig. 4(a), the particle decreases (increases) its tumbling period on the limit cycle when $T<T_{n}$ ($T>T_{n}$) by suitable shape deformations. For the tumbling dynamics, since the particle spends most of its time near the $\theta=0$ ghost region [38] (i.e. region of critical slowdown in state space), the shape modulations near this orientation have the most significant contribution to the tumbling period. This modulation is illustrated in Fig. 4(b) & (c) for $T<T_{n}$ and $T>T_{n}$ respectively. In these figures, the limit cycle is colored to denote the state-space speed $\sqrt{(dr/dt)^{2}+(d\theta/dt)^{2}}$. The limit cycle for the non-oscillatory system (red curve) is also plotted for comparison. We can see in Fig. 4(b) that compared to that of the non-oscillatory system, the oscillatory system reduces its shape parameter $r$ near the orientation $\theta=0$. The reduced elongation of the particle helps it traverse the ghost region faster and thus to reduce the period of the limit cycle. Similarly, the limit cycle in Fig. 4(c) has $r$ increased in order to traverse the ghost region slower and thus increase the period of the limit cycle. Hence, by appropriately modulating the shape near the ghost region, the particle matches the period of the limit cycle with the imposed oscillation period.
@@
Discussion and Outlook – We have shown that deformable anisotropic particles can phase-synchronize in time-dependent oscillatory shear flow. At the collective level, this results in the development of nematic order in a suspension of non-interacting particles. For $A\gg 1$, the mean shear flow component is superseded by the oscillatory component, and the particles undergo complex yet synchronized back-and-forth angular oscillations and tumbling motion (see Sec. ID-E of [36]). Our results also extend to continuum models of deformable particles by modifying $\beta(r)$, the flow aligning scale (see Sec. V of [36]). Furthermore, the phenomenon of synchronization-induced nematic order occurs in steady shear flow if the particles actively undergo shape oscillations or have periodically varying chiral activity (see Sec. III-IV of [36]). Such situations may be relevant for active particles undergoing shape deformations [39, 40, 41, 42] or in general when the active particles are undergoing periodic modulations in an internal state.

### v2

@@ Abstract
Unlike solids and simple liquids, orientational order of constituent particles is a defining feature of soft and biological matter [1, 2, 3, 4]. Classic theories predict the emergence of nematic order: Maier-Saupe theory based on anisotropic attractive interactions [5, 6] and Onsager’s theory based on the gain in translational entropy at the expense of lowered orientational entropy of the particles [7]. Athermal, active and external fields [8, 9, 10, 11, 12] can also drive the isotropic - nematic transition. Notably, all these approaches rely on interaction between anisotropic particles. They also ignore a characteristic feature ubiquitous in soft and biological matter, namely the deformability of the constituents e.g., as in polymers, soft colloids, emulsions and cells [13, 14]. In this letter, we show how shape deformability in conjunction with shape anisotropy can induce nematic order of non-interacting particles in a periodically driven system, via a phase synchronization mechanism.
@@
Figure 1: Emergence of nematic order for non-interacting deformable particles in an oscillatory shear flow. (a) Schematic of the system showing deformable particles suspended in a time-dependent shear flow with both a mean (red) and an oscillatory (blue) component. The initial isotropic state of random orientations $\theta_{i}(0)$ and varying shape parameters $r_{i}(0)$ develops into (b) a phase synchronized state at later times. Early and late time kymographs illustrating the temporal evolution of orientation of (c) deformable and (d) rigid particles. (e) The corresponding nematic order parameter $S(t)$ for deformable (black) and rigid (grey) particles. Grey ellipses at top of (c) are a guide for particle orientation while phase synchronized. The plots are for $N=200$, $A=0.5$ and $T=20$.
@@
Figure 3: Transient dynamics in state-space. Snapshots of trajectories in the state-space as points while passing through the $(r,2\theta)$ cross section of the torus at three different times. At $t=0$, random initial conditions (open circles); at $t=T$, trajectories converged to the limit cycle of the steady shear system (yellow filled circles on the red curve); and at $t=14T$, trajectories (black filled circles) converging onto the stable point (green filled circle with incoming arrows) that lies on the limit cycle of the oscillatory system (cyan curve). The unstable point (pink cross with outgoing arrows) corresponds to the unstable limit cycle in the full state-space.
Figure 4: (a) Arnold tongues: regions of phase synchronization induced nematic order (violet) in the $(A,T)$ parameter space. The inset shows the expanded view of the first tongue. (b)-(c) Limit cycles in $(r,2\theta)$ projection of the state-space, corresponding to two points marked in (a): $T=19.5\lesssim T_{n}$ in (b) and $T=20.7\gtrsim T_{n}$ in (c). The color on the limit cycle denotes the state-space speed $s=\sqrt{(dr/dt)^{2}+(d\theta/dt)^{2}}$. In each figure, the limit cycle for the case of steady shear ($T_{n}\approx 20.02$) is plotted in red. The grey and red ellipses illustrate the aspect ratio of the deformed particle (not to scale) as they pass through the ghost region (near the orientation $\theta=0$) in the oscillatory and in the steady shear flow respectively. Compared to particles in steady shear, in (b) particles reduce their anisotropy as they pass through the ghost region to reduce their period. Conversely, in (c) particles increase their anisotropy to increase their period.
@@
Emergence of nematic order – Simulations show that, over time, all particles tumble in the flow at the same driving frequency $\omega$, thus frequency locking their rotational motion to the imposed oscillatory shear flow. More importantly, we find that the particles also phase lock their rotational motion with the imposed flow (see Sec. IB of [37]). This phase locking of each particle’s orientation with the imposed oscillatory flow implies the synchronization in the motion of all particles with each other as well, and we obtain a state of collective behavior where orientations of all non-interacting deformable particles are the same at all, sufficiently long, times ($t\gg{2\pi}/{\omega}$). Thus, nematic order emerges in the system as indicated schematically in Fig. 1(b) (see Supplemental Video S1).
Kymographs illustrating the evolution of the orientation of particles $\theta_{i}(t)$ are shown in Fig. 1(c) for early and late times. The isotropy in the initial orientation of particles gradually fades away with time, and at long times the phase-synchronized state emerges with tumbling frequency $\omega$ and no variations in $\theta_{i}$ between different particles. In contrast, the kymograph obtained for rigid, elongated particles, shown in Fig. 1(d) does not exhibit phase locking and the system does not develop nematic order (see Supplemental Video S2).
@@
Phase synchronization – We now use a general mathematical framework [15] to demonstrate phase synchronization arising from particle deformability. Representing the dynamical system, Eqs. (1)-(2), in vector form,
$\dot{\mathbf{x}}_{i}=F(\mathbf{x}_{i};\dot{\gamma}),$
where $\mathbf{x}_{i}=(r_{i}(t),\theta_{i}(t))$, $F(\cdot)$ is the right-hand-side function, and $\dot{\gamma}(t)=1+A\sin(\omega t)$ is the parametric driving. For $A\ll 1$, doing a perturbation expansion up to linear order in $A$ we obtain
$\dot{\mathbf{x}}_{i}\approx F(\mathbf{x}_{i})|_{\dot{\gamma}=1}+A\,\frac{\partial F(\mathbf{x}_{i})}{\partial\dot{\gamma}}\Bigg{|}_{\dot{\gamma}=1}\sin(\omega t).$
Performing a phase reduction [15] from the state variables $(r,\theta)$ to a phase variable $\psi$ that varies along the oscillation (see Sec. II of [37]), and integrating the resulting equation over one period of the parametric driving $T=2\pi/\omega$, we obtain the circle map
$\psi_{k+1}=\psi_{k}+2\pi\frac{\omega_{n}}{\omega}+A\,f(\psi_{k}),$ (3)
[…]
Dynamical analysis– Eqs. (1)-(2) constitute a non-autonomous dynamical system due to the time-dependent driving $\dot{\gamma}(t)$, but introducing a new phase variable $\phi_{i}=\phi=\omega t$ or,
@@
with $\phi(t+2\pi/\omega)=\phi(t)$, results in Eqs. (1), (2) and (4) to form an autonomous $3$D dynamical system for a single particle  (see Sec. IA of [37] for more details) and allow us to pursue a state-space analysis.
@@
Physical mechanism– Figure 4(a) shows the range of amplitude, $A$, and oscillatory driving period, $T$, for which phase synchronization occurs. The synchronization regions (violet) occur periodically in bands near multiples of $T_{n}$, the natural time period of tumbling in a steady shear flow ($A=0$). These structures in Fig. 4(a) are the well-known Arnold tongues [15] and they come in and out of existence in parameter space via saddle-node bifurcation of cycles in state space (see Sec. IC of [37]).
@@
The finite width of the tongues can be physically rationalized by analyzing the extent to which a deformable particle can modulate the rate of tumbling and shape changes to match the driving period $T$ of the oscillatory shear. For example, within the primary Arnold tongue in Fig. 4(a), the particle decreases (increases) its tumbling period on the limit cycle when $T<T_{n}$ ($T>T_{n}$) by suitable shape deformations. For the tumbling dynamics, since the particle spends most of its time near the $\theta=0$ ghost region [40] (i.e. region of critical slowdown in state space), the shape modulations near this orientation have the most significant contribution to the tumbling period. This modulation is illustrated in Fig. 4(b) & (c) for $T<T_{n}$ and $T>T_{n}$ respectively. In these figures, the limit cycle is colored to denote the state-space speed $s=\sqrt{(dr/dt)^{2}+(d\theta/dt)^{2}}$. The limit cycle for the non-oscillatory system (red curve) is also plotted for comparison. We can see in Fig. 4(b) that compared to that of the non-oscillatory system, the oscillatory system reduces its shape parameter $r$ near the orientation $\theta=0$. The reduced elongation of the particle helps it traverse the ghost region faster and thus to reduce the period of the limit cycle. Similarly, the limit cycle in Fig. 4(c) has $r$ increased in order to traverse the ghost region slower and thus increase the period of the limit cycle. Hence, by appropriately modulating the shape near the ghost region, the particle matches the period of the limit cycle with the imposed oscillation period.
@@
Discussion and Outlook – We have shown that deformable anisotropic particles can phase-synchronize in time-dependent oscillatory shear flow. At the collective level, this results in the development of nematic order in a suspension of non-interacting particles. For $A\gg 1$, the mean shear flow component is superseded by the oscillatory component, and the particles undergo complex yet synchronized back-and-forth angular oscillations and tumbling motion (see Sec. IE-IF of [37]). Our results also extend to continuum models of deformable particles by modifying $\beta(r)$, the flow aligning scale (see Sec. V of [37]). Furthermore, the phenomenon of synchronization-induced nematic order occurs in steady shear flow if the non-interacting particles are active. In Sec. III and IV of [37], we provide mathematical analysis and numerical results to show phase synchronization for particles actively undergo shape oscillations or have periodically varying chiral activity, respectively. Such situations may be relevant for active particles undergoing shape deformations [41, 42, 43, 44] or in general when the active particles are undergoing periodic modulations in an internal state. Hence, our work opens up new avenues to explore at the intersection of nonlinear dynamics of synchronization and soft & active matter.